第4章離散確率変数
4.1 Probability Distribution Function (PDF) for a Discrete Random Variable4.1 離散確率変数の確率分布関数(PDF)
A discrete probability distribution function has two characteristics:
離散確率分布関数には、二つの特徴がある。
- Each probability is between zero and one, inclusive.
英語のヒント
各確率は0以上1以下である。 - The sum of the probabilities is one.確率の合計は1である。
Example 4.1例題 4.1
A child psychologist is interested in the number of times a newborn baby's crying wakes its mother after midnight. For a random sample of 50 mothers, the following information was obtained. Let X = the number of times per week a newborn baby's crying wakes its mother after midnight. For this example, x = 0, 1, 2, 3, 4, 5.
ある児童心理学者は、新生児の泣き声で母親が深夜0時以降に目を覚ます回数に関心をもっている。母親50人の無作為標本から、次の情報が得られた。X = 新生児の泣き声で母親が深夜0時以降に目を覚ます1週間当たりの回数とする。この例では、x = 0, 1, 2, 3, 4, 5である。
P(x) = probability that X takes on a value x.
P(x) = Xが値xを取る確率。
x |
P(x) |
|---|---|
0 |
P(x = 0) = |
1 |
P(x = 1) = |
2 |
P(x = 2) = |
3 |
P(x = 3) = |
4 |
P(x = 4) = |
5 |
P(x = 5) = |
X takes on the values 0, 1, 2, 3, 4, 5. This is a discrete PDF because:
Xは0、1、2、3、4、5の値を取る。これは次の理由から離散PDFである。
- Each P(x) is between zero and one, inclusive.各P(x)は0以上1以下である。
- The sum of the probabilities is one, that is,確率の合計は1である。すなわち、
Example 4.2例題 4.2
Suppose Nancy has classes three days a week. She attends classes three days a week 80% of the time, two days 15% of the time, one day 4% of the time, and no days 1% of the time. Suppose one week is randomly selected.
ナンシーには週3日の授業があるとする。週3日出席する割合は80%、週2日は15%、週1日は4%、全く出席しない割合は1%である。1週間を無作為に選ぶとする。
a. Let X = the number of days Nancy ____________________.
a. X = ナンシーが____________________日数とする。
b. X takes on what values?
b. Xはどのような値を取るか。
c. Suppose one week is randomly chosen. Construct a probability distribution table (called a PDF table) like the one in Example 4.1. The table should have two columns labeled x and P(x). What does the P(x) column sum to?
c. 1週間を無作為に選ぶとする。例4.1のような確率分布表(PDF表)を作成せよ。表にはxとP(x)という見出しの2列を設ける。P(x)の列の合計はいくらになるか。
Solution解き方
a. Let X = the number of days Nancy attends class per week.
a. X = ナンシーが1週間に授業に出席する日数とする。
b. 0, 1, 2, and 3
b. 0、1、2、3
c.
x |
P(x) |
|---|---|
0 |
0.01 |
1 |
0.04 |
2 |
0.15 |
3 |
0.80 |
Chapter Review章の復習
The characteristics of a probability distribution function (PDF) for a discrete random variable are as follows:
離散確率変数の確率分布関数(PDF)の特徴は、次のとおりである。
- Each probability is between zero and one, inclusive (inclusive means to include zero and one).各確率は0以上1以下である(両端の0と1も含む)。
- The sum of the probabilities is one.確率の合計は1である。
Exercises 1–5練習問題 1–5
Use the following information to answer the next five exercises: A company wants to evaluate its attrition rate, in other words, how long new hires stay with the company. Over the years, they have established the following probability distribution.
英語のヒント
次の情報を使って、続く5問に答えよ。ある会社が離職率、つまり新規採用者がどれほど長く会社にとどまるかを評価しようとしている。長年の調査から、次の確率分布を得た。
Let X = the number of years a new hire will stay with the company.
X = 新規採用者が会社にとどまる年数とする。
Let P(x) = the probability that a new hire will stay with the company x years.
P(x) = 新規採用者が会社にx年間とどまる確率とする。
Exercise 1問題 1
x |
P(x) |
|---|---|
0 |
0.12 |
1 |
0.18 |
2 |
0.30 |
3 |
0.15 |
4 |
|
5 |
0.10 |
6 |
0.05 |
Solution解き方
x |
P(x) |
|---|---|
0 |
0.12 |
1 |
0.18 |
2 |
0.30 |
3 |
0.15 |
4 |
0.10 |
5 |
0.10 |
6 |
0.05 |
Exercise 2問題 2
P(x = 4) = _______
Exercise 3問題 3
P(x ≥ 5) = _______
Solution解き方
0.10 + 0.05 = 0.15
Exercise 4問題 4
On average, how long would you expect a new hire to stay with the company?
平均して、新規採用者はどれほど長く会社にとどまると予想されるか。
Exercise 5問題 5
What does the column “P(x)” sum to?
「P(x)」の列の合計はいくらになるか。
Solution解き方
1
Exercises 6–9練習問題 6–9
Use the following information to answer the next six exercises: A baker is deciding how many batches of muffins to make to sell in his bakery. He wants to make enough to sell every one and no fewer. Through observation, the baker has established a probability distribution.
英語のヒント
次の情報を使って、続く6問に答えよ。あるパン職人が、店で売るマフィンを何回分焼くか決めようとしている。作ったものがすべて売れ、しかも不足しない量を作りたいと考えている。観察から、次の確率分布を得た。
x |
P(x) |
|---|---|
1 |
0.15 |
2 |
0.35 |
3 |
0.40 |
4 |
0.10 |
Exercise 6問題 6
Define the random variable X.
確率変数Xを定義せよ。
Exercise 7問題 7
What is the probability the baker will sell more than one batch? P(x > 1) = _______
パン職人が1回分より多く売る確率はいくらか。P(x > 1) = _______
Solution解き方
0.35 + 0.40 + 0.10 = 0.85
Exercise 8問題 8
What is the probability the baker will sell exactly one batch? P(x = 1) = _______
パン職人がちょうど1回分を売る確率はいくらか。P(x = 1) = _______
Exercise 9問題 9
On average, how many batches should the baker make?
平均して、パン職人は何回分作るべきか。
Solution解き方
1(0.15) + 2(0.35) + 3(0.40) + 4(0.10) = 0.15 + 0.70 + 1.20 + 0.40 = 2.45
Exercises 10–12練習問題 10–12
Use the following information to answer the next four exercises: Ellen has music practice three days a week. She practices for all of the three days 85% of the time, two days 8% of the time, one day 4% of the time, and no days 3% of the time. One week is selected at random.
次の情報を使って、続く4問に答えよ。エレンには週3日の音楽の練習がある。3日とも練習する割合は85%、2日は8%、1日は4%、全く練習しない割合は3%である。1週間を無作為に選ぶ。
Exercise 10問題 10
Define the random variable X.
確率変数Xを定義せよ。
Exercise 11問題 11
Construct a probability distribution table for the data.
このデータの確率分布表を作成せよ。
Solution解き方
x |
P(x) |
|---|---|
0 |
0.03 |
1 |
0.04 |
2 |
0.08 |
3 |
0.85 |
Exercise 12問題 12
We know that for a probability distribution function to be discrete, it must have two characteristics. One is that the sum of the probabilities is one. What is the other characteristic?
離散確率分布関数には二つの特徴が必要である。一つは確率の合計が1であることだ。もう一つの特徴は何か。
Exercises 13–17練習問題 13–17
Use the following information to answer the next five exercises: Javier volunteers in community events each month. He does not do more than five events in a month. He attends exactly five events 35% of the time, four events 25% of the time, three events 20% of the time, two events 10% of the time, one event 5% of the time, and no events 5% of the time.
次の情報を使って、続く5問に答えよ。ハビエルは毎月、地域行事でボランティアをしている。1か月に5件を超える行事には参加しない。ちょうど5件に参加する割合は35%、4件は25%、3件は20%、2件は10%、1件は5%、全く参加しない割合は5%である。
Exercise 13問題 13
Define the random variable X.
確率変数Xを定義せよ。
Solution解き方
Let X = the number of events Javier volunteers for each month.
X = ハビエルが毎月ボランティアとして参加する行事の数とする。
Exercise 14問題 14
What values does x take on?
xはどのような値を取るか。
Exercise 15問題 15
Construct a PDF table.
PDF表を作成せよ。
Solution解き方
x |
P(x) |
|---|---|
0 |
0.05 |
1 |
0.05 |
2 |
0.10 |
3 |
0.20 |
4 |
0.25 |
5 |
0.35 |
Exercise 16問題 16
Find the probability that Javier volunteers for less than three events each month. P(x < 3) = _______
ハビエルが毎月ボランティアとして参加する行事が3件未満である確率を求めよ。P(x < 3) = _______
Exercise 17問題 17
Find the probability that Javier volunteers for at least one event each month. P(x > 0) = _______
ハビエルが毎月少なくとも1件の行事にボランティアとして参加する確率を求めよ。P(x > 0) = _______
Solution解き方
1 – 0.05 = 0.95